n-3/4 + n+5/7= 5/4
Answer by stanbon(75887)

n-3/4 + n+5/7= 5/4
---------------
Multiply thru by 28 to get:
7(n-3) + 4(n+5) = 7*5
-------
7n-21+4n+20 = 35
-------
11n-1 = 35
11n = 36
-----
n = 3 3/11
-----------------
Cheers,
Stan H.